OpenStudy (anonymous):

Given the equation P2 = A3, what is the orbital period, in years, for the planet Saturn? (Saturn is located 9.5 AU from the sun.) 5 years 29 years 229 years 500 years

OpenStudy (anonymous):

P^2 = period squared = A^3 = AU distance cubed. Find A^3 then take its square root to get P.